Brad is 7 more than 2 times as old as Ben. In 23 years the sum of their ages will be 83. How old is Ben now?
I have been trying to figure out a way to set up the problem to work it out.
Brad = y 2x + 7+ 23y = 83
Ben = x x + 23y = 83
I don't want the answer, I just need help setting it up to work out.

Answer by mananth(16946)

Let ben’s age be x
So Brad’s age will be 2x+7
After 23 years their ages will be
x+23 and (2x+7)+23
x+23 +(2x+7)+23=83
x+23 + 2x+7+23= 83
3x+53=83
3x=30
X=10 --------------- Ben’s age
Enjoy your maths